# Order Out of Noise and Disorder: Fate of the Frustrated Manifold

### v1

@@ 3.9 Comparison: Manifold-Dependent Dimension Reduction
Table 13: Dynamic dimension reduction across manifolds (dimensionless units). The reduction is approximate due to thermal and interaction-induced fluctuations. Statistics are from full time series analysis ($t=0$ to $50$) for $N=400$ particles at $T=0.4$ with Gaussian couplings. Steady-state values shown where applicable.
@@
$\langle\Theta\rangle$ (steady) 9.6-9.6 $45.8$ $80.5$
$\sigma_{\Theta}$ (steady) $148.7$ $243.8$ $396.0$
@@
Figure 9 provides a unified view of rotational dynamics across all three geometries, enabling direct comparison of the symmetry-breaking phenomena.
Figure 9: Comparison of rotational dynamics across geometries. Langevin dynamics with random interactions: $N=400$, $T=0.4$. Top row: Total cumulative rotation time series. (a) Sphere $S^{2}$: Total $\Theta_{z}$ with steady-state $\langle\Theta_{z}\rangle=-9.6$, $\sigma=148.7$. (b) Cylinder: Total $\Theta_{\theta}$ with steady-state $\langle\Theta_{\theta}\rangle=45.8$, $\sigma=243.8$. (c) Torus $T^{2}$: Total $\Theta_{\theta}$ with steady-state $\langle\Theta_{\theta}\rangle=80.5$, $\sigma=396.0$. Bottom row: Per-structure angular velocities. (d) Sphere: Band $\omega_{z}$ showing coherent oscillations. (e) Cylinder: Per-cluster $\omega_{\theta}$ for four clusters (C0–C3), showing independent fluctuations. (f) Torus: Per-ring $\omega_{\theta}$ for two rings, showing weak anti-correlation ($\rho=-0.27$). The comparison reveals progressively richer correlation structure: single coherent band (sphere) $\to$ four weakly anti-correlated clusters (cylinder) $\to$ two weakly anti-correlated rings (torus).
While dimension reduction from 2D to quasi-1D structures occurs on all three manifolds, the character of the reduction depends strongly on the specific manifold. The shared features are:
  1. 1.
    2D $\to$ 1D collapse: All three manifolds exhibit collapse from the full 2D surface to quasi-1D structures (circles, localized clusters, or disjoint unions thereof).
  2. 2.
    Approximate reduction: The collapse is never perfect; residual width and fluctuations persist due to thermal noise and interaction disorder.
  3. 3.
    Glassy relaxation: The approach to the low-dimensional attractor proceeds through multiple timescales characteristic of glassy systems.
However, the manifold-specific differences are equally notable, reflecting the interplay of topology (global connectivity) and local geometry (metric structure, curvature):
  • On the sphere, the uniform local geometry (constant positive Gaussian curvature) and simple topology produce the cleanest behavior. The Langevin equation (67) predicts isotropic diffusion of $\mathbf{L}$ with effective coefficient $D_{L}=2ND/3$ for uniform distributions. The single equatorial band exhibits coherent dynamics with diffusive rotation (ensemble-averaged MSD exponent $\alpha\approx 0.9$), directly confirming the pure-diffusion prediction with no deterministic drift. Individual realizations show $\alpha$ ranging from $0.55$ to $1.13$ due to quenched disorder effects, but ensemble averaging recovers the theoretical $\alpha=1$.
  • On the bounded cylinder, the presence of boundaries (a topological feature) dramatically alters the dynamics. The finite axial direction $[0,H]$ discretizes into two levels, while interaction frustration further fragments the azimuthal distribution into localized clusters. The four-cluster configuration exhibits weak anti-correlations between cluster angular velocities ($\rho\approx-0.1$ to 0.2-0.2). The cylinder shows ballistic behavior with ensemble-averaged MSD exponent $\alpha\approx 2.0$, robust across five independent realizations. This contrasts with the theoretical prediction of pure diffusion, suggesting that the boundary conditions create persistent collective modes that do not decorrelate. The clusters, pinned at reflective $z$-boundaries, may sustain long-lived rotational excitations that maintain coherence over the simulation timescale.
  • On the torus, the product topology $S^{1}\times S^{1}$ (closed, like the sphere, but with different fundamental group) combines with non-uniform local geometry (position-dependent metric factor $(R+r\cos\varphi)^{2}$) to produce qualitatively distinct behavior. Rings at different poloidal positions evolve largely independently, with different effective moment arms $(R+r\cos\varphi)$ contributing to $\Theta_{\theta}$. This induces weak anti-correlation between ring angular velocities ($\rho=-0.27$). The torus exhibits superdiffusive behavior with ensemble-averaged MSD exponent $\alpha\approx 1.7$, intermediate between the diffusive sphere and ballistic cylinder. Notably, all five realizations show systematic positive drift ($\langle\bar{\omega}\rangle=+0.075$, $p<0.001$), which arises from a stochastic ratchet mechanism: the asymmetric ring positions combined with position-dependent noise amplitude create a non-equilibrium steady state that rectifies thermal fluctuations into directed rotation.
@@ 3.10 Symmetry Analysis: Steady States and Disorder-Induced Rotational Symmetry Breaking
Despite the apparent similarity, our phenomenon differs fundamentally from spontaneous symmetry breaking (SSB) as understood in quantum field theory [13]. In true SSB, the ground state does not share the Hamiltonian’s symmetry, the symmetry-breaking direction is fixed once selected, and Goldstone modes emerge corresponding to motion along the degenerate ground-state manifold. In our system, the symmetry-breaking direction (band orientation, cluster positions, ring locations) is not fixed but slowly evolves in time, driven by thermal noise. The theoretical framework of Section 2.4 makes this precise: the cumulative rotation $\Theta$ performs a random walk according to the stochastic differential equations in Table 1, with geometry-dependent effective diffusion coefficients. Pairwise forces produce zero net torque (equation (23) and its analogues for other geometries), so these equations contain no deterministic drift, only pure diffusion. True SSB also requires the thermodynamic limit $N\to\infty$; our finite-$N$ structures are precursors to any putative SSB.
Goldstone Modes and Soft Collective Excitations.
When spacetime symmetries (such as e.g. rotations) are broken, the counting of Goldstone modes changes dramatically from the internal symmetry case. Two distinct classification schemes exist in the literature, corresponding to Hamiltonian and dissipative systems respectively:
  • Nielsen-Chadha classification (Hamiltonian systems) [23, 24]: Goldstone modes are classified as Type-I (linear dispersion $\omega\propto k$) or Type-II (quadratic dispersion $\omega\propto k^{2}$). The number of modes may be fewer than the number of broken generators. Physical examples include: ferromagnets (one Type-II magnon with $\omega\propto k^{2}$); antiferromagnets (two Type-I magnons with $\omega\propto k$).
  • Minami-Hidaka classification (dissipative systems) [2, 3]: When continuous symmetry is spontaneously broken in an open system coupled to a heat bath (as in Langevin dynamics), the gapless Goldstone modes are generically diffusive rather than propagating. These are classified as type-A diffusion modes (with dispersion $\omega=-iDk^{2}$, i.e., relaxation rate $\Gamma\propto k^{2}$) or type-B diffusion modes (with $\omega=-iD^{\prime}k^{4}$ or gapped), depending on whether the Hamiltonian Noether charge $Q_{R}$ is conserved. When the heat bath breaks conservation of $Q_{R}$, as in Langevin dynamics, type-A modes become diffusive.
The correspondence between these classifications is: Hamiltonian Type-II modes (quadratic dispersion) become type-A diffusion modes in dissipative systems. Since our model is a dissipative Langevin system, the Minami-Hidaka classification applies, and we expect type-A diffusion modes.
Our slowly drifting collective variables constitute a concrete realization of these dissipative type-A Goldstone modes. The connection to the theoretical framework emerges through the following observations:
@@
    Diffusive MSD on the sphere: Our simulations show $\langle\Delta\theta^{2}\rangle\propto t^{\alpha}$ with ensemble-averaged $\alpha\approx 0.9$ on the sphere, confirming pure diffusive scaling. This is the real-space signature of type-A diffusion modes at $k=0$. However, the torus ($\alpha\approx 1.7$) and cylinder ($\alpha\approx 2.0$) deviate significantly from diffusion.
@@
    Power spectrum: For a diffusive process, the Wiener-Khinchin theorem relates the MSD to the power spectrum. A random walk with $\langle\Delta\theta^{2}\rangle\propto t$ implies $S_{\theta}(\omega)\propto 1/\omega^{2}$ at low frequencies. This $1/\omega^{2}$ spectrum is observed in our cumulative rotation time series (Figure 7a).
  3. 3.
    Zero mode of gapless spectrum: In the Minami-Hidaka framework, type-A diffusion modes have relaxation rate $\Gamma\propto Dk^{2}$. Our $q=0$ (uniform rotation) mode has $\Gamma=0$, the zero mode of a gapless excitation spectrum, consistent with the type-A diffusion classification of [3].
  4. 4.
    Geometry-dependent dynamics: A key finding is the geometry dependence of MSD scaling: the sphere shows $\alpha\approx 1$ (diffusive), confirming theoretical predictions for type-A Goldstone modes, while the cylinder ($\alpha\approx 2$, ballistic) and torus ($\alpha\approx 1.7$, superdiffusive with systematic drift) deviate significantly. Interestingly, all three geometries exhibit the $1/\omega^{2}$ power spectrum despite different MSD exponents, suggesting the spectrum reflects local Langevin dynamics while the MSD is sensitive to global constraints from boundaries and metric structure.
Table 15: Summary of MSD scaling and power spectrum behavior across geometries. Only the sphere exhibits the predicted diffusive MSD ($\alpha\approx 1$) from type-A Goldstone mode theory. The torus shows superdiffusive behavior ($\alpha\approx 1.7$) with systematic drift, while the cylinder is ballistic ($\alpha\approx 2.0$). Interestingly, all geometries share the $1/\omega^{2}$ power spectrum despite different MSD exponents.
@@
Geometry Residual Sym. MSD $\alpha$ $S_{\Theta}(\omega)$ scaling Goldstone manifold
Sphere $S^{2}$ SO(2) (continuous) $0.9$ (diffusive) $\propto 1/\omega^{2}$ $S^{2}\cong\text{SO}(3)/\text{SO}(2)$
[…]
Cylinder $\mathbb{Z}_{2}$ (discrete) $2.0$ (ballistic) $\propto 1/\omega^{2}$ None (persistent modes)
The MSD scaling reveals a clear geometry dependence: the sphere exhibits diffusive behavior ($\alpha\approx 1$), while the cylinder shows ballistic scaling ($\alpha\approx 2$) and the torus is superdiffusive ($\alpha\approx 1.7$). These results are robust under ensemble averaging over multiple independent realizations.
The diffusive behavior on the sphere directly confirms the theoretical prediction of pure rotational diffusion with no deterministic drift; a consequence of the zero-torque property of the Langevin dynamics. However, the deviations on the cylinder and torus indicate that boundary conditions and metric structure can qualitatively modify the collective dynamics even when the underlying Langevin equation predicts diffusion.
On the sphere specifically, the band orientation lives on $S^{2}\cong\text{SO}(3)/\text{SO}(2)$ and performs a random walk on this Goldstone manifold. The theoretical prediction (67) gives the rotational diffusion coefficient $D_{\text{rot}}=2ND/3$ for uniformly distributed particles, implying $\langle\Delta\theta^{2}\rangle\propto D_{\text{rot}}t$. The ensemble-averaged MSD exponent $\alpha=0.88\pm 0.26$ confirms this prediction: individual realizations show substantial variation ($\alpha=0.55$ to $1.13$) due to quenched disorder effects, but ensemble averaging recovers the theoretical diffusive scaling.
The cylinder and torus, by contrast, deviate significantly from pure diffusion. The cylinder’s ballistic behavior ($\alpha\approx 2$) indicates that the four-cluster configuration sustains persistent collective modes that maintain rotational coherence over the simulation timescale. The boundary conditions may prevent efficient relaxation of collective rotation, creating long-lived rotational excitations. The torus’s superdiffusive behavior ($\alpha\approx 1.7$) combined with systematic positive drift ($\langle\bar{\omega}\rangle=+0.075$, consistent across all five realizations) arises from a stochastic ratchet mechanism. When rings form at asymmetric poloidal positions ($\varphi\approx 295^{\circ}$ and $\varphi\approx 125^{\circ}$), they experience different effective noise amplitudes $\sigma_{\theta}(\varphi)\propto 1/(R+r\cos\varphi)$. Thermal fluctuations in $\varphi$ around these mean positions are rectified by the nonlinear coupling between radial position and azimuthal diffusivity, producing net rotation, a geometry-induced Brownian motor effect.
@@ 4.1 Connections to Spin Glass Physics
The band precession on the sphere, correlated cluster-pair oscillations on the cylinder, and dual ring rotations on the torus represent slow collective modes that continue to evolve long after dimensional reduction is complete. These modes are analogous to the “soft modes” in structural glasses. Full time series analysis ($t=0$ to $50$) reveals manifold-dependent behavior: on the sphere, the band forms by $t\approx 5$ with diffusive rotation ($\alpha=0.9\pm 0.3$); on the cylinder, four clusters form rapidly by $t\approx 1$ with ballistic rotation ($\alpha=2.0\pm 0.05$); on the torus, two rings form by $t\approx 13$ with superdiffusive rotation ($\alpha=1.7\pm 0.1$) and systematic positive drift ($\langle\bar{\omega}\rangle=+0.075$).

### v2

@@ 3.9 Comparison: Manifold-Dependent Dimension Reduction
Table 12: Dynamic dimension reduction across manifolds (dimensionless units). Statistics are from full time series analysis ($t=0$ to $50$) for $N=400$ particles at $T=0.4$ with Gaussian couplings.
@@
While dimension reduction from 2D to quasi-1D structures occurs on all three manifolds, the character of the reduction depends strongly on the specific manifold. The shared features are: (i) 2D $\to$ 1D collapse to quasi-1D structures; (ii) approximate reduction with residual width and fluctuations from thermal noise; and (iii) glassy relaxation through multiple timescales.
The manifold-specific differences reflect the interplay of topology and local geometry:
  • On the sphere, uniform local geometry and simple topology produce the cleanest behavior: diffusive rotation ($\alpha\approx 0.9$) with no systematic drift, directly confirming the pure-diffusion prediction.
  • On the bounded cylinder, boundaries dramatically alter the dynamics. The four-cluster configuration shows ballistic behavior ($\alpha\approx 2.0$), suggesting boundary conditions create persistent collective modes that do not decorrelate.
  • On the torus, non-uniform local geometry produces superdiffusive behavior ($\alpha\approx 1.7$) with systematic positive drift arising from a stochastic ratchet mechanism.
@@
3.10 Summary: Orientation Dynamics Across Geometries
Table 13 summarizes the structure orientation statistics for all three geometries.
Table 13: Structure orientation statistics across geometries. The formation phase shows larger thickness (diffuse structure) and larger orientation fluctuations. In NESS, structures are well-formed (small $\mathcal{T}$) with characteristic orientation dynamics. Simulations with $N=400$ particles, $T=0.4$, transition time $t^{*}=13$.
Geometry Phase Thickness $\mathcal{T}$ $\boldsymbol{\sigma(n_{z})}$ Correlation time $\tau_{c}$
Sphere $S^{2}$ Formation $0.62\pm 0.15$ $0.19$
NESS $0.29\pm 0.12$ $0.11$ $18.4$
Cylinder Formation $36\pm 0.11 0.0$90.09
NESS $34\pm 0.09 0.0$70.07 $3.2$
Torus $T^{2}$ Formation $0.27\pm 0.07$ $0.32$
NESS $0.27\pm 0.08$ $0.44$ $18.4$
The structure normal analysis reveals a clear hierarchy of orientation dynamics:
  1. 1.
    Sphere: Full 3D precession on $S^{2}$ with isotropic exploration, long correlation time ($\tau_{c}\approx 18$), and diffusive dynamics. The band normal is a true order parameter living on the Goldstone manifold $\mathrm{SO}(3)/\mathrm{SO}(2)\cong S^{2}$.
  2. 2.
    Torus: Constrained precession with $z$-bias, reflecting the intrinsic geometry of ring configurations on $T^{2}$. Long correlation time comparable to the sphere, but with larger amplitude fluctuations due to the discrete (two-ring) structure.
  3. 3.
    Cylinder: Planar precession confined to the $xy$-plane, with short correlation time ($\tau_{c}\approx 3.2$) reflecting rapid reorganization of the discrete cluster configuration. The boundary-induced localization constrains orientation dynamics to a 1D subspace of $S^{2}$.
Connection to Dissipative Nambu-Goldstone Modes.
The structure normal $\hat{\mathbf{n}}(t)$ is the Goldstone mode variable: it specifies where the emergent structure is oriented on the Goldstone manifold $S^{2}$ (for the sphere) or related spaces (for other geometries). The autocorrelation function $C(\tau)=\langle\hat{\mathbf{n}}(t)\cdot\hat{\mathbf{n}}(t+\tau)\rangle$ and its Fourier transform, the structure normal power spectrum $S_{n_{z}}(\omega)$, characterize this Goldstone mode dynamics. For type-A diffusion modes in dissipative systems, theory predicts $S(\omega)\propto 1/\omega^{2}$ at low frequencies [5, 6]. Our measurements reveal geometry-dependent power law exponents: $S_{n_{z}}(\omega)\propto\omega^{-\beta}$ with $\beta\approx 1.6$ on the sphere, $\beta\approx 2.0$ on the cylinder, and $\beta\approx 1.4$ on the torus. Only the cylinder approaches the theoretical $1/\omega^{2}$ prediction, but the cylinder has no continuous Goldstone manifold (SO(2) $\to\mathbb{Z}_{2}$). The deviations on the sphere and torus may reflect finite-size effects or the bounded nature of $\hat{\mathbf{n}}$ on $S^{2}$.
The long correlation time $\tau_{c}\approx 18$ on the sphere and torus is the characteristic timescale for the Goldstone mode, far exceeding the microscopic thermal correlation time ($\tau_{v}<0.2$). This separation of timescales confirms the adiabatic nature of orientation dynamics: the structure can be treated as quasi-static on timescales shorter than $\tau_{c}$. The short correlation time on the cylinder ($\tau_{c}\approx 3.2$) reflects the absence of a continuous Goldstone manifold; the discrete $\mathbb{Z}_{2}$ symmetry allows rapid reorientation between equivalent configurations.
@@ 3.11 Symmetry Analysis: Steady States and Disorder-Induced Rotational Symmetry Breaking
Despite the apparent similarity, our phenomenon differs fundamentally from spontaneous symmetry breaking (SSB) as understood in quantum field theory [16]. In true SSB, the ground state does not share the Hamiltonian’s symmetry, the symmetry-breaking direction is fixed once selected, and Goldstone modes emerge corresponding to motion along the degenerate ground-state manifold. In our system, the symmetry-breaking direction (band orientation, cluster positions, ring locations) is not fixed but slowly evolves in time, driven by thermal noise. Pairwise forces produce zero net torque (equation (23) and its analogues for other geometries), so there is no deterministic restoring force that would freeze the orientation. True SSB also requires the thermodynamic limit $N\to\infty$; our finite-$N$ structures are precursors to any putative SSB.
Identifying the Goldstone Mode: Structure Normal vs. Cumulative Rotation.
A key conceptual point requires clarification. In our system, two distinct observables characterize rotational dynamics:
@@
    Structure normal $\hat{\mathbf{n}}(t)$: The slow variable that specifies where the emergent structure is oriented in 3D space. This is the Goldstone mode variable that explores the manifold of degenerate configurations (the Goldstone manifold).
@@
    Cumulative rotation $\Theta(t)$: The fast variable that measures total angular displacement around a fixed axis. This is not a Goldstone mode; it is driven by the instantaneous angular momentum, which thermalizes rapidly in the overdamped limit.
The analogy with ferromagnetism makes this distinction clear. In a ferromagnet, SO(3) $\to$ SO(2) breaking produces one Goldstone mode: the slow reorientation of the magnetization direction $\hat{\mathbf{M}}(t)$ on $S^{2}\cong\text{SO}(3)/\text{SO}(2)$. The total spin $|\mathbf{M}|$ is fixed; only its direction fluctuates. Similarly, on our sphere, the band normal $\hat{\mathbf{n}}(t)$ plays the role of $\hat{\mathbf{M}}$, slowly precessing on the Goldstone manifold $S^{2}$. The cumulative rotation $\Theta_{z}(t)$ measures something different: how much the system has rotated around the external $z$-axis, regardless of band orientation. This is analogous to measuring total angular momentum transfer, not magnetization direction.
Goldstone Mode Counting.
The number of Goldstone modes depends on the symmetry breaking pattern:
  • Sphere: SO(3) $\to$ SO(2) breaks two generators. The Goldstone manifold is $S^{2}$, a 2-dimensional space. In dissipative systems, this yields one diffusive mode: the slow precession of $\hat{\mathbf{n}}(t)$ on $S^{2}$, with correlation time $\tau_{c}\approx 18$.
  • Torus: SO(2) $\times$ SO(2) $\to$ SO(2) $\times\mathbb{Z}_{2}$ breaks one continuous generator (toroidal SO(2) $\to\mathbb{Z}_{2}$). The Goldstone manifold is $S^{1}/\mathbb{Z}_{2}$, corresponding to the collective toroidal position of the ring pair. The structure normal $\hat{\mathbf{n}}(t)$ captures a related but distinct aspect: the orientation of the ring plane, which has a preferred direction due to torus geometry ($\langle n_{z}\rangle\approx 0.73$).
  • Cylinder: SO(2) $\to\mathbb{Z}_{2}$ breaks continuous symmetry to discrete. There is no continuous Goldstone manifold; the discrete $\mathbb{Z}_{2}$ symmetry allows only finite rotations ($180^{\circ}$) between equivalent configurations. This explains the short correlation time ($\tau_{c}\approx 3.2$) and ballistic MSD: the system does not explore a continuous manifold but instead exhibits persistent rotational modes.
Dissipative Goldstone Mode Classification.
When spacetime symmetries are broken in dissipative systems, the Minami-Hidaka classification [5, 6] applies:
  • Type-A diffusion modes: Relaxation rate $\Gamma\propto Dk^{2}$. At $k=0$, this gives a gapless diffusive mode. The power spectrum prediction is $S(\omega)\propto 1/\omega^{2}$ at low frequencies.
  • Type-B diffusion modes: Relaxation rate $\Gamma\propto D^{\prime}k^{4}$ or gapped.
For our sphere, the slow precession of $\hat{\mathbf{n}}(t)$ should exhibit type-A behavior. Our measurements of the structure normal power spectrum $S_{n_{z}}(\omega)\propto\omega^{-\beta}$ yield:
  • Sphere: $\beta\approx 1.6$, somewhat shallower than the $1/\omega^{2}$ prediction
  • Cylinder: $\beta\approx 2.0$, consistent with $1/\omega^{2}$ (but this geometry has no continuous Goldstone manifold)
  • Torus: $\beta\approx 1.4$, the shallowest spectrum
The deviations from $\beta=2$ on the sphere and torus may reflect: (i) finite-size effects; (ii) the bounded nature of $\hat{\mathbf{n}}$ on $S^{2}$, which constrains long-time fluctuations; (iii) coupling between different components of $\hat{\mathbf{n}}$; or (iv) the non-uniform metric on the torus creating additional correlations.
The Fast Cumulative Rotation.
The cumulative rotation $\Theta(t)$ exhibits different behavior that should not be interpreted as Goldstone physics:
  • On the sphere, $\Theta_{z}(t)$ shows diffusive MSD ($\alpha\approx 0.9$), consistent with the fast, thermally-driven rotational dynamics predicted by the overdamped Langevin equation.
  • On the cylinder, $\Theta_{z}(t)$ shows ballistic MSD ($\alpha\approx 2.0$), indicating persistent collective modes that do not decorrelate.
  • On the torus, $\Theta_{\theta}(t)$ shows superdiffusive MSD ($\alpha\approx 1.7$) with systematic drift from the stochastic ratchet mechanism.
The diffusive scaling $\alpha\approx 1$ on the sphere confirms that $\Theta(t)$ is driven by white thermal noise, as expected for the fast angular momentum variable in the overdamped limit. This is the behavior of a driven system, not the slow exploration of a Goldstone manifold.
Table 15: Summary of Goldstone mode characteristics across geometries. The structure normal $\hat{\mathbf{n}}(t)$ represents the Goldstone mode (slow orientation dynamics), while the cumulative rotation $\Theta(t)$ represents fast angular momentum dynamics. The sphere provides the cleanest Goldstone behavior; the cylinder has no continuous Goldstone manifold; the torus has geometric complications.
@@
Geometry Goldstone manifold $\tau_{c}$ of $\hat{\mathbf{n}}$ $S_{n_{z}}\propto\omega^{-\beta}$ MSD $\alpha$ of $\Theta$
Sphere $S^{2}$ $S^{2}$ (continuous) $18.4$ $\beta\approx 1.6$ $0.9$ (diffusive)
Torus $T^{2}$ $S^{1}/\mathbb{Z}_{2}$ (w/ drift) $18.4$ $\beta\approx 1.4$ $1.7$ (superdiffusive)
Cylinder None (discrete) $3.2$ $\beta\approx 2.0$ $2.0$ (ballistic)
@@ 4.1 Connections to Spin Glass Physics
The band precession on the sphere, correlated cluster-pair oscillations on the cylinder, and dual ring rotations on the torus represent slow collective modes that continue to evolve long after dimensional reduction is complete. These modes are analogous to the “soft modes” in structural glasses. Full time series analysis ($t=0$ to $50$) reveals manifold-dependent behavior: on the sphere, the band forms by $t\approx 5$ with diffusive rotation ($\alpha=0.9\pm 0.3$); on the cylinder, four clusters form rapidly by $t\approx 1$ with ballistic rotation ($\alpha=2.0\pm 0.05$); on the torus, two rings form by $t\approx 13$ with superdiffusive rotation ($\alpha=1.7\pm 0.1$) and systematic positive drift ($\langle\bar{\omega}\rangle=+0.075$). The structure orientation analysis of Section 3.10 reveals that these slow modes have correlation times $\tau_{c}\approx 18$ on the sphere and torus, but only $\tau_{c}\approx 3$ on the cylinder, reflecting the different “stiffness” of continuous versus discrete structures against thermal reorientation.
@@
In mean-field spin glasses, the Parisi order parameter distribution $P(q)$ characterizes replica overlap and exhibits non-trivial structure below the glass transition [1, 37]. The cumulative rotation $\Theta(t)$ in our model is fundamentally different: it is a time-integrated quantity that grows diffusively, $\langle\Theta^{2}\rangle\propto D_{\text{rot}}t$, rather than a bounded order parameter. At any fixed time $t$, the distribution $P(\Theta;t)$ is approximately Gaussian (from the central limit theorem applied to the sum of many particle contributions), with zero ensemble mean (from rotational symmetry) and variance that grows linearly with time.